solve ethe linear system by using substitution:
6x-y=-35
5x-2y=-35

Answer by benazir.sj@gmail.com(109)

6x-y=-35-------------(1)
5x-2y=-35-------------(2)
since both the equations are equal to -35
(1)=(2)
6x-y =5x-2y
6x-5x=y-2y
x=-y
now substitute the value of x in any one of the equation
x=-y
6x-y= -35
6(-y) -y=-35
-7y = -35
y=-35/-7
y=5
put y = 5 in (1)
6x-y =-35
6x-5=-35
6x =-35+5
6x=-30
x= -30/6
x=-5
ie,
x= -5
and y= 5
Proof:
6x-y=-35
6(-5) -5 =-35
-30-5=-35
-35=-35
or
5x-2y =-35
5(-5) -2*5 =-35
-25-10 =-35
-35=-35
LHS =RHS